### v1

@@ 5.1 Refracted Brownian risk model
From Theorem 4.5 and Proposition 5.1, we have $c_{2}^{*}\in[\varepsilon_{2},\infty)$, and with Proposition 4.2, $c_{1}^{*}\in[0,\varepsilon_{2}]$. Accordingly, the two-dimensional optimization problem to determine the values of $c_{1}^{*}$ and $c_{2}^{*}$ can be addressed by solving two auxiliary one-dimensional optimization problems. Specifically, $c_{1}^{*}$, which is the unique minimizer of the function $H(\cdot,c_{2})$ on the interval $0,\varepsilon_{2}]$, is the unique intersection point of $\vartheta^{(q+m,q)}{{}^{\prime}}(\cdot,-l)$ and $H(\cdot,c_{2})$. Then, $c_{2}^{*}$ is determined as the unique point in $\varepsilon_{2},\infty)$ such that $\vartheta^{(q+m,q)}{{}^{\prime}}(c_{2}^{*},-l)=\vartheta^{(q+m,q)}{{}^{\prime}}(c_{1}^{*},-l)$. Based on this, we can directly derive the following result, for more details, see also Section 4 in Loeffen[27].
@@ 5.2 Refracted Cramér-Lundberg model
where $r_{2}=\frac{\sqrt{(\mu\alpha-q-\eta)^{2}+4\mu q\alpha}-\mu\alpha+q+\eta}{2\mu}$, $r_{1}=\frac{-\sqrt{(\mu\alpha-q-\eta)^{2}+4\mu q\alpha}-\mu\alpha+q+\eta}{2\mu}$, $G_{i}=\frac{\alpha+r_{i}}{r},\ (i=1,2)$ and $r=r_{2}-r_{1}=\frac{\sqrt{(\mu\alpha-q-\eta)^{2}+4\mu q\alpha}}{\mu}$. Similar to the one above, superscripts * and $Y$ are for $W^{(q+m)}$ and $\mathbb{W}^{(q)}$ respectively. From $\frac{r_{i}}{r_{i}-r_{2}^{Y}}=-\frac{\mu-\delta}{\delta}\frac{r_{i}-r_{1}^{Y}}{r_{i}+\alpha}$ and $\frac{G_{2}^{Y}r_{i}}{r_{i}-r_{2}^{Y}}-\frac{G_{1}^{Y}r_{i}}{r_{i}-r_{1}^{Y}}=-\frac{\mu-\delta}{\delta},\ (i=1,2)$ it follows that, for $x\geq b$,
@@ Appendix A Proofs and more
In the following we provide proofs of propositions and lemmas for completeness.
@@
A.3 Proof of Lemma 3.3

### v2

@@ 5.1 Refracted Brownian risk model
From Theorem 4.6 and Proposition 5.1, we have $c_{2}^{*}\in[\varepsilon_{2},\infty)$, and with Proposition 4.2, $c_{1}^{*}\in[0,\varepsilon_{2}]$. Accordingly, the two-dimensional optimization problem to determine the values of $c_{1}^{*}$ and $c_{2}^{*}$ can be addressed by solving two auxiliary one-dimensional optimization problems. The numerical procedure is as follows.
Remark 5.2.
For given model parameters, we first pre-compute $\vartheta^{(q+m,q)}(x,-l)$ and its derivative on a fine uniform grid. Proposition 5.1 yields constants $\varepsilon_{1}\leq 0\leq\varepsilon_{2}$ at which the monotonicity of $\vartheta^{(q+m,q)\prime}(\cdot,-l)$ changes. By Theorem 4.6 and Proposition 5.1 we have $c_{2}^{*}\in[\varepsilon_{2},\infty)$, and Proposition 4.2 guarantees that, for each fixed $c_{2}>\varepsilon_{2}$, the function $H(\cdot,c_{2})$ admits a unique minimiser $c_{1}^{*}(c_{2})$ on $0,\varepsilon_{2})$, obtained as the unique intersection of $\vartheta^{(q+m,q)}{{}^{\prime}}(\cdot,-l)$ and $H(\cdot,c_{2})$ on this interval. Finally, $c_{2}^{*}$ is determined as the unique solution of $\vartheta^{(q+m,q)}{{}^{\prime}}(c_{2},-l)=\vartheta^{(q+m,q)}{{}^{\prime}}(c_{1}^{*}(c_{2}),-l)$ on $\varepsilon_{2},\bar{c}_{2}]$ for a sufficiently large upper bound $\bar{c}_{2}$, solved numerically as a one-dimensional root-finding problem.
Alternatively, restricting $c_{1}\in[0,\varepsilon_{2}]$ and $c_{2}\in[\varepsilon_{2},\infty)$, the optimal pair $(c_{1}^{*},c_{2}^{*})$ can also be identified as an intersection of the level sets $\vartheta^{(q+m,q)\prime}(c_{1},-l)=H(c_{1},c_{2})$ and $\vartheta^{(q+m,q)\prime}(c_{2},-l)=\vartheta^{(q+m,q)\prime}(c_{1},-l)$, located numerically via contour plots. If no intersection arises in this region, the optimal pair is selected among the boundary configurations in cases (ii)-(v) of Proposition 4.2, by comparing the associated values of $H$.
Based on this, we can directly derive the following result, for more details, see also Section 4 in Loeffen[27].
@@ 5.2 Refracted Cramér-Lundberg model
where $r_{2}=\frac{\sqrt{(\mu\alpha-q-\eta)^{2}+4\mu q\alpha}-\mu\alpha+q+\eta}{2\mu}$, $r_{1}=\frac{-\sqrt{(\mu\alpha-q-\eta)^{2}+4\mu q\alpha}-\mu\alpha+q+\eta}{2\mu}$, $G_{i}=\frac{\alpha+r_{i}}{r},\ (i=1,2)$ and $r=r_{2}-r_{1}=\frac{\sqrt{(\mu\alpha-q-\eta)^{2}+4\mu q\alpha}}{\mu}$. Similar to the one above, superscripts * and $Y$ are for $W^{(q+m)}$ and $\mathbb{W}^{(q)}$ respectively. For $i\in\{1,2\}$ we write $\i\{1,2\}\backslash$ i for the index in $\{1,2\}$ different from $i$. From $\frac{r_{i}}{r_{i}-r_{2}^{Y}}=-\frac{\mu-\delta}{\delta}\frac{r_{i}-r_{1}^{Y}}{r_{i}+\alpha}$ and $\frac{G_{2}^{Y}r_{i}}{r_{i}-r_{2}^{Y}}-\frac{G_{1}^{Y}r_{i}}{r_{i}-r_{1}^{Y}}=-\frac{\mu-\delta}{\delta},\ (i=1,2)$ it follows that, for $x\geq b$,
@@ Appendix A Proofs and more
In the following we provide proofs of propositions, corollaries and lemmas for completeness.
@@
A.3 Proof of Corollary 3.2
Proof.
By (2.1), (2.2) and (2.9), we have
$\vartheta^{(q+m,q)}(x,-l)$
$=W^{(q+m)}(x+l)-m\int_{0}^{x}W^{(q)}(x-y)W^{(q+m)}(l+y)\mathrm{d}y+\delta\mathbf{1}{\{x\geq b\}}\int_{b}^{x}\mathbb{W}^{(q)}(x-y)$
$\ \ \ \cdot\Big(W^{(q+m)}{{}^{\prime}}(y+l)-m\big(W^{(q)}(0)W^{(q+m)}(y+l)+\int_{0}^{y}W^{(q)}{{}^{\prime}}(y-z)W^{(q+m)}(z+l)\mathrm{d}z\big)\Big)\mathrm{d}y,$
since
$W^{(q)}{{}^{\prime}}(y+l)+m\int_{0}^{l}W^{(q)}{{}^{\prime}}(y+l-z)W^{(q+m)}(z)\mathrm{d}z$
$=\frac{\mathrm{d}}{\mathrm{d}y}\bigl(W^{(q+m)}(y+l)-m\int_{0}^{y}W^{(q)}(y-z)W^{(q+m)}(z+l)\mathrm{d}z\bigr)$
$=W^{(q+m)}{{}^{\prime}}(y+l)-m\bigl(W^{(q)}(0)W^{(q+m)}(y+l)+\int_{0}^{y}W^{(q)}{{}^{\prime}}(y-z)W^{(q+m)}(z+l)\mathrm{d}z\bigr).$
From Kyprianou[17] it follows that $\lim_{l\to\infty}\tfrac{W^{(q)}(x+l)}{W^{(q)}(l)}=e^{\phi(q)x}$ and $\lim_{l\to\infty}$ $\tfrac{{W^{(q)}}{{}^{\prime}}(x+l)}{W^{(q)}(l)}=\phi(q)e^{\phi(q)x}$. Dividing both the numerator and the denominator in the above expression by $W^{(q+m)}(l)$ and then letting $l\to\infty$, we obtain (3.2). ∎
A.4 Proof of Lemma 3.6